Eunice is 18 years younger than her father. if her father is three times her current age, how old are they?
Found 2 solutions by mila nedic, MathTherapy:
Answer by mila nedic(19)

Let's say that Eunice's age will be represented by the letter e, and her father's age will be represented by the letter f. Let's make an equation using the information we have.
If she is 18 years younger than her father, that mean we can use the variables e and f to create an equation: e+18=f
The second piece of information we are given is that her father is 3 times her current age. We can also use the variables to make an equation of with this information: 3e=f
e+18=f
3e=f
Now we see that both equation are actually equal because they are both equal to f so to find e we can use that. We can write it as e+18=3e and now we can solve it.
e+18=3e
18=2e (we subtracted e from each side so we could get all variables on one side)
e=9 (divided each side by 18 to isolate e)
Now we know that Eunice is 9 years old and we can use either one of the starting equations 3e=f and e+18=f to figure out her father's age, by inputting 9 instead of e since we figured that out already.
e+18=f
9+18=f
27=f (her father is 24, but we can use the other equation too to make sure that the answers are right)
3e=f
3*9=f
27=f (since we got the same answer, we know that we got it right)
Therefore Eunice is 9 years old and her father is 27 years old.

Answer by MathTherapy(10556)


Eunice is 18 years younger than her father. if her father is three times her current age, how old are they?
Let Eunice's age be E
Then her dad's is: 3E
We then get: 3E - E = 18
2E = 18
Eunice, or
Now, you should have no problems finding her dad's age!